OpenStudy (anonymous):

3 Square root of 5 multiplied by 2 with the square root of 4 equals

OpenStudy (anonymous):

You don't understand I want to know how to work the problem. I don't get it!

OpenStudy (anonymous):

9

OpenStudy (anonymous):

Where did the other 3 come from?

OpenStudy (kc_kennylau):

\[3\sqrt5\times2\sqrt4=3\sqrt5\times2\times2=12\sqrt5\]

OpenStudy (anonymous):

Okay I see now thank you

OpenStudy (kc_kennylau):

Note that it is not "3 to the power 5" @javo22

OpenStudy (kc_kennylau):

and @leahtripplethree I still have no idea what the hell "3 Square root of 5" means.

OpenStudy (kc_kennylau):

Therefore I am just guessing.

OpenStudy (whpalmer4):

"3 square root of 5" if written \[3\sqrt{5} = 3 * \sqrt{5}\] Now, if you saw this: \[\sqrt[3]{5}\]that means the 3rd (or cube) root of 5: a number that when multiplied with itself twice = 5. For example, \[\sqrt[3]{8} = 2\]because \[2*2*2 = 8\]

OpenStudy (kc_kennylau):

Therefore the question is ambiguous.

OpenStudy (kc_kennylau):

However I reckon that it is \(3\sqrt5\) since one would not write \(\sqrt[2]4\).

OpenStudy (whpalmer4):

no, the problem isn't ambiguous if we see it accurately. Hearing your description of what it looks like on the paper, however, is a different story.

OpenStudy (whpalmer4):

Without knowing whether you are aware of how a cube root might be notated, we can't really be sure which of the two is meant, but I agree that \(3\sqrt{5}\) is more likely because you get an answer that reflects actual useful manipulation, unlike with the cube root where all that is reflected is the ability to figure out the square root of 4 :-)

OpenStudy (whpalmer4):

sorry, @kc_kennylau I said "your description" but was referring to the original poster's description, of course!

OpenStudy (kc_kennylau):

lol